Consider the HF solution with the concentration of 5%(w/v). Calculate its concentration in the molarity.
Answer #1

The definition of a 5% w/v solution is: 5g of solute dissolved in 100mL of solution

Now you know that you have 5g of HF dissolved in water to make a final solution volume of 100mL
That is equivalent to 50g HF in 1000mL of solution or 1.0L of solution
Molar mass HF = 20.0g/mol
Mol of HF = 50/20 = 2.5 mol HF dissolved in 1.0L solution
Molarity of HF solution = 2.5M
Answer #2

The definition of a 5% w/v solution is: 5g of solute dissolved in 100mL of solution

Now you know that you have 5g of HF dissolved in water to make a final solution volume of 100mL
That is equivalent to 50g HF in 1000mL of solution or 1.0L of solution
Molar mass HF = 20.0g/mol
Mol of HF = 50/20 = 2.5 mol HF dissolved in 1.0L solution
Molarity of HF solution = 2.5M
Answer #3

Now you know that you have 5g of HF dissolved in water to make a final solution volume of 100mL
That is equivalent to 50g HF in 1000mL of solution or 1.0L of solution
Molar mass HF = 20.0g/mol
Mol of HF = 50/20 = 2.5 mol HF dissolved in 1.0L solution
Molarity of HF solution = 2.5M (answer)
Answer #4

let volume of solution=100 mL

mass of solute=mass of HF=5 g

mole of HF=mass/molar mass=5/20=0.25 mole

molarity=mole/volume=0.25/100*10-3=2.5 M